Assume that the number of days needed to hatch an egg of a certain type of...

  1. Assume that the number of days needed to hatch an egg of a certain type of a rare lizard is distributed normally. Using an incubator, we were able to hatch 13 eggs from different nests. We have a sample mean of 18.97 weeks with a sample standard variance of 10 weeks. What is the 90% confidence interval for the population variance?
Answer #1

Solution :  

Given that,

c = 90%

s2 = 10

n = 13

d.f. = n - 1 = 13 -1 = 12

At 90% confidence level ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.05 and 1 - (\alpha / 2) = 0.95

Now , using chi square table ,

\chi ^{2}\alpha /2,df = \chi ^{2} 0.05,12 = 21.026

\chi ^{2}1-\alpha /2,df = \chi ^{2} 0.95,12 = 5.226

The 90% confidence interval for \sigma 2 is,

(n-1)s Xa/2,df < 2 (n-1)s Xi-a/2),df 2

(13 - 1) *10 / 21.026 < \sigma 2 < (13 -1 ) *10 /5.226

5.707 < \sigma 2 < 22.962

The 90% confidence interval for \sigma 2 is  (5.707 , 22.962)